allison invested $2200 in two banks. at the first bank she had an interest rate of 4 1/2 % and at the second bank the interest rate was 5 3/4%. if the interest obtained at the end of the year was $107.75,how much money was invested at 5 3/4% answer:$700 were invested at 5 3/4 % need a break down of how they got $700 were invested at 5 3/4%
Answer by checkley77(12844)

.0575x+.045(2200-x)=107.75
.0575x+99-.045x=107.75
.0125x=107.75-99
.0125x=875
x=875/.0125
x=700 Amount invested @ 5.75%
2200-700=1,500 Amount invested @ 4.5%
Proof:
.0575*700+.045*1500=107.75
40.25+67.5=107.75
107.75=107.75